The work function of a certain metal is φ = 3.05 eV. Determine the minimum frequency of light f0 for which photoelectrons are emitted from the metal. (Planck's constant is: h = 4.1357×10-15 eVs.)

Determine the corresponding wavelength of light. (Possibly useful constant: hc = 1240 eVnm.)

Calculate the kinetic energy of the emitted electrons if the frequency of the photons is doubled to 2f0.
Answer #1

\varphi =hf_o\\\\f_o=\frac{\varphi }{h}=\frac{3.05}{4.1357\times 10^{-15}}=7.37\times 10^{14}Hz\\\\\\wavelength\; \lambda \;= \frac{c}{f} = \frac{3\;x\;10^8}{7.37\;x\;10^{14}}= 4.07\;x\;10^{-7} m\;=\;407\;nm


===================================

kinetic \;energy \;K \;= \;h\;f - h\;f_o\\\\ K = h\;2\;f_o - h\;f_o\\\\\\ K = hf_o = 3.05 eV